The height of the Washington monument is 50.5 inches more than 10 times the length of a side of its square base. If the sum of these two dimensions is 7327 inches, find the height of the Washington monument.
Answer by edjones(8007)

h=10L+50.5
h+L=7327
10L+50.5+L=7327
11L=7276.5
L=661.5"
h=6615+50.5=6665.5"
.
Ed